18 July 1999 Owl shift summary

Valeri Jejer (jejer@fnal.gov)
Sun, 18 Jul 1999 07:05:47 -0500 (CDT)

Shift Summary for owl shift, 18/Jul/99

On Shift: Theo, Valeri

Hours of beam available at all intensities: 7.00
Hours of beam available at nominal intensity: 7.00
Hours of E832 data (special runs not included): 7.00
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:

~10 min - kfrends

========================================================================
Runs ended this shift:
Run 13941 - in progress - BEAM_832

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13941 71-73 00:57 Ktevana Errors - missing FFFF word in Cr.53,
kfrend.

141 02:25 Laser stab. RMS wide, no action taken.

147-148 02:33 DYC-53 busy - kfrend.

168 03:02 DYC-53 busy - kfrend.

03:05 Filling gas checklist noticed that He bag 1
pressure is fluctuating between -.034 and 0
within a minute or two.

172 03:07 Ktevana Errors - unknown crates in stream 5,
kfrend.

228 04:20 DYC-39 busy - kfrend.

243 04:40 Ktevana Errors - unknown crates in stream 5,
kfrend.

05:15 Onsplit requests QKN013 in tps3d2, done.

287 05:40 Ktevana Errors - kfrend.

06:00 Called MCR asking to increase intensity to 4.2e12
(were getting 4e12).

325 06:32 Laser stab. RMS wide, no action taken.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

AZLAT in crate 53, slot 13 should be replaced - page Jason H. during next
significant non-owl downtime.